Flowering Plants

An index of common and scientific names

This is an index to the Chart of Flowering Plant Families (for availability, see below). All names appearing on the chart are included in the index. In addition, we include many further scientific names and also common names of useful and ornamental plants and plant products, thus giving a more extensive listing than is possible on the chart itself.

The index may be used to locate common names, families and other names on the chart using the hierarchical structure as follows. Given a common name or plant product, for example Mango, the entry in the index says that Mango is in the genus Mangifera, in the family Anacardiaceae. This family has an entry giving the order as Sapindales, which is in the Malvids, in the Rosids, which, finally, is in the Eudicots, one of the major taxa of flowering plants.

Warning: The common, or vernacular, names in this index cover familiar plants and plant products. These names vary from language to language. Moreover, often only local names are available. Sometimes, the same name is used for several (sometimes quite unrelated) species. We have tried to disambiguate common names using genera. However, some common names can be used quite loosely of a range of different species (usage sometimes varying around the world). Particular problems arise with names of timbers and also with some descriptive names of garden flowers. Where ambiguity would result, we have omitted such names. In all, the common names here should be viewed as an indicator of family and genus for many specific plants and products and as a guide to current usage of some common names.

Resources

For a comprehensive index of flowering plants and ferns consult:

Mabberley’s Plant-book: A Portable Dictionary of Plants, their Classifications, and Uses, 3rd Edition, David J. Mabberley (2008).

A valuable resource is the website:

Angiosperm Phylogeny Website. Version 9, June 2008 [and more or less continuously updated since], P.F. Stevens, (2001 onwards). URL http://www.mobot.org/MOBOT/research/APweb/

Families in this Index are (an updated version of) those of:

An update of the Angiosperm Phylogeny Group classification for the orders and families of flowering plants: APG III. Angiosperm Phylogeny Group. 2009. Botanical Journal of the Linnean Society, 2009, 161:105-121.

Availability

This document and the Chart of Flowering Plant Families are available from the Botanical Chart website at http://www.cs.manchester.ac.uk/flora.

Index

Abelia, Caprifoliaceae Abolbodaceae, in Xyridaceae Abutilon, Malvaceae Acacia, Fabaceae Acai palm (Euterpe), Arecaceae Acanthaceae, Lamiales Acanthus, Acanthaceae Aceraceae, in Sapindaceae Acerola (Malpighia), Malpighiaceae Achariaceae, Malpighiales Achatocarpaceae, Caryophyllales Achiote (Bixa), Bixaceae Acoraceae, Monocots Actinidiaceae, Ericales Adoxaceae, Dipsacales Aextoxicaceae, Berberidopsidales Afara (Terminalia), Combretaceae Afina wood (Strombosia), Strombosiaceae African blackwood (Dalbergia), Fabaceae African dogrose (Oncoba), Salicaceae African dogrose (Xylotheca), Achariaceae African forest elder (Nuxia), Stilbaceae African hemp (Sparrmannia), Malvaceae African oak (Lophira), Ochnaceae African oak (Oldfieldia), Picrodendraceae African violet (Saintpaulia), Gesneriaceae African walnut (Coula), Coulaceae African white pear (Apodytes),

Icacinaceae (provisionally) Afrostyrax, Huaceae Agapanthus, Amaryllidaceae Agapanthaceae, in Amaryllidaceae Agar wood (Aquilaria & sometimes Gyrinops),

Thymelaeaceae Agavaceae, in Asparagaceae Agave, Asparagaceae Agdestis, Phytolaccaceae Ageratum, Asteraceae Airplants (Tillandsia), Bromeliaceae Aizoaceae, Caryophyllales Akaniaceae, Brassicales Akebia, Lardizabalaceae Akee (Blighia), Sapindaceae Alangiaceae, in Cornaceae Alder (Alnus), Betulaceae Alfalfa (Medicago), Fabaceae Alismataceae, Alismatales Alismatales, Monocots Alkanet (various genera, including Anchusa

& Alkanna), Boraginaceae Alliaceae, in Amaryllidaceae Allspice (Pimenta), Myrtaceae Allspice, Carolina (Calycanthus), Calycanthaceae Almond (Prunus), Roseae Aloe, Xanthorrhoeaceae Alseuosmiaceae, Asterales Alstroemeriaceae, Liliales Altingiaceae, Saxifragales Alyssum, Brassicaceae Alzateaceae, Myrtales Amaranthaceae, Caryophyllales Amaryllidaceae, Asparagales Amborellaceae, Amborellales Amborellales, Early-branching order Ambrette (Abelmoschus), Malvaceae Amelanchier, Rosaceae Amphorogynaceae, in Santalaceae Anacardiaceae, Sapindales Anarthriaceae, Poales Ancistrocladaceae, Caryophyllales Andaman marblewood (Diospyros), Ebenaceae Androsace, Primulaceae Aneulophus, Erythroxylaceae Angelica, Apiaceae Angostura (bark) (Galipea

& Cusparia = Angostura), Rutaceae Angueuk wood (Ongokea), Aptandraceae Anemone, Ranunculaceae Anise (Pimpinella), Apiaceae Aniseed myrtle (Syzygium), Myrtaceae Anisophyllaceae, Cucurbitales Annatto (Bixa), Bixaceae Annonaceae, Magnoliales Anthurium, Araceae Aphelandra, Acanthaceae Aphyllanthaceae, in Asparagaceae Apiaceae, Apiales Apiales, Campanulides Apocynaceae, Gentianales Apodanthaceae, Cucurbitales (provisionally) Apodytes, Icacinaceae (provisionally) Aponogetonaceae, Alismatales Apostasiaceae, in Orchidaceae Apple (Malus), Rosaceae Apple, Custard (Annona), Annonaceae Apple, Kei (Dovyalis = Aberoa), Salicaceae Apple, Star (Chrysophyllum), Sapotaceae Apricot (Prunus), Rosaceae Aptandraceae, Santalales Aquifoliaceae, Aquifoliales Aquifoliales, Campanulids Araceae, Alismatales Araliaceae, Apiales Aralidiaceae, in Torricelliaceae Ardisia, Primulaceae Areca, Arecaceae Arecaceae, Arecales Arecales, Monocots Argan oil (Argania), Sapotaceae Argophyllaceae, Asterales Aristolochiaceae, Piperales Arracacha (Arracacia), Apiaceae Arrowgrass, various genera in Juncaginaceae

and Scheuchzeriaceae Arrowhead (Sagittaria), Alismataceae Arrowhead (Nephthytis & Syngonium), Araceae Arrowroot, various genera and common names,

including: ‘East Indian arrowroot’ (Curcuma), Zingiberaceae ‘Polynesian arrowroot’ (Tacca), Dioscoreaceae ‘West Indian arrowroot’ (Maranta), Marantaceae Artemisia, Asteraceae Artichoke, Globe (Cynara), Asteraceae Artichoke, Jerusalem (Helianthus), Asteraceae Arum, Araceae Asclepiadaceae, in Apocynaceae Asafoetida (Ferula), Apiaceae Ash (Fraxinus), Oleaceae Ash, Silver (Flindersia), Rutaceae Ashoka tree (Saraca), Fabaceae Asparagus, Asparagaceae Asparagaceae, Asparagales Asparagales, Moncots Aspen (Populus), Salicaceae Asphodel, (Asphodelus), Xanthorrhoeaceae Asphodel, Bog (Narthecium), Nartheciaceae Asphodelaceae, in Xanthorrhoeaceae Aspidistra, Asparagaceae Astelia, Asteliaceae Asteliaceae, Asparagales Asteraceae, Asterales Asterales, Campanulids Asterids, Eudicots Astilbe, Saxifragaceae Athel tree (Tamarix), Tamaricaceae Atherospermataceae, Laurales Aubergine (Solanum), Solanaceae Aucuba, Garryaceae Australian flycatcher (Cephalotus), Cephalotaceae Australian heath (Epacris), Ericaceae Austrobaileyaceae, Austrobaileyales Austrobaileyales, early-branching order Averrhoaceae, in Oxalidaceae Avicennia, Acanthaceae Avicenniaceae, in Acanthaceae Avocado (Persea), Lauraceae Azalea (Rhododendron), Ericales

Bael (Aegle), Rutaceae Baku (Tieghemella), Sapotaceae Balanopaceae, Malpighales Balanophoraceae, Santalales Balsa (Ochroma), Malvaceae Balsam, Copaiba (Copaifera), Fabaceae Balsam of Mecca (Commiphora), Burseraceae Balsam, Peru (Myroxylon), Fabaceae Balsam poplar (Populus), Salicaceae Balsam, Tolu (Myroxylon), Fabaceae Bacuri (Platonia), Clusiaceae Balsa (Ochroma), Malvaceae Balsam (Impatiens), Balsaminaceae Balsam of Peru (Myroxylon), Fabaceae Balsaminaceae, Ericales Bamboo (various genera), Poaceae Bamboo, Sacred/Heavenly (Nandina), Berberidaceae Bamburanta (Ctenanthe), Marantaceae Bambusaceae, in Poaceae Banana (Musa & Ensete), Musaceae Banksia, Proteaceae Banyan (Ficus), Moraceae Baobab (Adansonia), Malvaceae Barbados cherry (Malpighia), Malpighiaceae Barberry (Berberis), Berberidaceae Barbeuiaceae, Caryopyllales Barbeyaceae, Rosales Barclaya, Nymphaeaceae Barclayaceae, in Nymphaeaceae Barley (Hordeum), Poaceae Barringtonia, Lecythidaceae Barringtoniaceae, in Lecythidaceae Basellaceae, Caryophyllales Basil (Ocimum), Lamiaceae Basswood [USA] (Tilia), Malvaceae Bastard bullet tree (Humiria), Humiriaceae Bat flower (Tacca), Dioscoreaceae Bataceae, Brassicales Batidaceae = Bataceae Baueraceae, in Cunoniaceae Bay tree/leaves (Laurus), Lauraceae Bay, Loblolly (Gordonia), Theaceae Bay, West Indian (Pimenta), Myrtaceae Bayberry (Myrica), Myricaceae Bdellium (Commiphora), Burseraceae Bean (Phaseolus, Vigna, Vicia etc), Fabaceae Beautyberry (Callicarpa), Lamiaceae Bedstraw (Galium), Rubiaceae Bee balm (Monarda), Lamiaceae Beebrush (Aloysia), Verbenaceae Beech (Fagus), Fagaceae Beech, Southern (Nothofagus), Nothofagaceae Beech, White (Gmelina), Lamiaceae Beech myrtle (Nothofagus), Nothofagaceae Beefsteak plant (Perilla), Lamiaceae Beet (Beta), Amaranthaceae Begonia, Begoniaceae Begoniaceae, Cucurbitales Belladonna (Atropa), Solanaceae Bellflowers (Campanula, Campanulastrum, Codonopsis, Platycodon, etc.), Campanulaceae Ben nut (Moringa), Moringaceae Bengal quince (Aegle), Rutaceae Beni-seed (Sesame), Pedaliaceae Benzoin resin (Styrax), Styracaceae Berberidaceae, Ranunculales Bergamot (Monarda), Lamiaceae Bergamot orange (Citrus), Rutaceae Berzelia, Bruniaceae Betel pepper (Piper), Piperaceae Betony (Stachys), Lamiaceae Betulaceae, Fagales Bhesa, Centroplacaceae Biebersteiniaceae, Sapindales Bignoniaceae, Lamiales Bilberry (Vaccinium), Ericaceae Bilimbi (Averrhoa), Oxalidaceae Bindweed (Convolvulus & Calystegia),

Convolvulaceae Bindweed, Black (Fallopia), Polygonaceae Bindweed, Blue (Solanum), Solanaceae Birch (Betula), Betulaceae Bird-of-Paradise (Strelitzia), Strelitziaceae Birthwort (Aristolochia), Aristolochiaceae Bischofiaceae, in Phyllanthaceae Bittersweet (Solanum), Solanaceae Bittersweet (Celastrus), Celastraceae Bixaceae, Malvales Black mangrove, Black (Avicenna), Acanthaceae Black-eyed Susie/Susan (Hibiscus), Malvaceae Black-eyed Susie/Susan (Rudbeckia), Astraceae Black-eyed Susie/Susan (Thunbergia), Acanthaceae Blackberry (Rubus), Rosaceae Bladder-flower (Araujia), Apocynaceae Bladder nut (Staphylea), Staphyleaceae Bladderwort (Utricularia), Lentibulariaceae Blandfordiaceae, Asparagales Blazing star (Liatris), Asteraceae, Blazing star (Mentzelia), Loasaceae Blepharocaryaceae, in Anacardiaceae Blinks (Montia), Montiaceae Bloodwort (Achillea), Asteraceae Bloodwort (Rumex), Polygonaceae Bloodwort/Bloodroot (Sanguinaria), Papaveraceae Bloodworts (various genera), Haemodoraceae Bloodberry (Cordia), Boraginaceae, Bloodberry (Rivina), Phytolaccaceae Blueberry (Vaccinium), Ericaceae Bo tree (Ficus), Moraceae Bog asphodel (Narthecium), Nartheciaceae Bog myrtle (Myrica), Myricaceae Bog-rosemary (Andromeda), Ericaceae Bogbean (Menyanthes), Menyanthaceae Bogota tea (Symplocos), Symplocaceae Bomarea, Alstroemeriaceae Bombacaceae, in Malvaceae Bonnetiaceae, Malpighiales Boobialla (Myoporum), Scrophulariaceae Boojum tree (Fouquieria), Fouquieriaceae Borage (Borago), Boraginaceae Boraginaceae, Lamiids Boryaceae, Asparagales Bottlebrush (Beaufortia), Myrtaceae Bottlebrush (Callistemon), Myrtaceae Bottlebrush, Natal (Greyia), Melianthaceae Bougainvillea, Nyctaginaceae Box (Buxus), Buxaceae Boxwood, West Indian (Casearia), Salicaceae Boxwood, West Indian (Phyllostylon), Ulmaceae Boxwood, West Indian (Tabebuia), Bignoniaceae Bramble (Rubus), Rosaceae Brassicaceae, Brassicales Brassicales, Malvids Brazil nut (Bertholletia), Lecythidaceae Brazil wax (Copernicia), Arecaceae Brazilian tulipwood (Dalbergia), Fabaceae Brazilwood or Pau-Brasil (Caesalpinia), Fabaceae Breadfruit (Artocarpus), Moraceae Breadfruit, Nicobar (Pandanus), Pandanaceae Breadnut (Brosimum and Artocarpus), Moraceae Bretschneidera, Akaniaceae Bromeliaceae, Poales Broom (Cytisus, Genista & Chamaecytisus), Fabaceae Broomrape (Orobanche), Orobanchaceae Brunelliaceae, Oxalidales Brunia, Bruniaceae Bruniaceae, Bruniales Bruniales, Campanulids Brunoniaceae, in Goodeniaceae Brushholly (Xylosma), Salicaceae Bryony, White (Bryonia), Cucurbitaceae Bryony, Black (Tamus = Dioscorea), Dioscoreaceae Buckeye [USA] (Aesculus), Sapindaceae Buckthorn (Rhamnus), Rhamnaceae Buckthorn, Sea (Hippopha), Elaeagnaceae Buckwheat (Fagopyrum), Polygonaceae Buckwheat, Wild (Eriogonum & Fallopia),

Polygonaceae Buddleia = Buddleja, Scrophulariaceae Buddlejaceae, in Scrophulariaceae Bugle (Ajuga), Lamiaceae Bulrush, various genera in Cyperaceae, also [UK] Typha, Typhaceae Bur-reed (Sparganium), Sparganiaceae Burma, Pride of/Jewel of (Curcuma), Zingiberaceae Burmanniaceae, Dioscoreales Burseraceae, Sapindales Bush mango (Irvingia), Irvingiaceae Bushmint (Hyptis & Mentha), Lamiaceae Busy Lizzie/Lizzy (Impatiens), Balsaminaceae Butcher’s broom (Ruscus), Asparagaceae Butomaceae, Alismatales Buttercup (Ranunculus), Ranunculaceae Butterfly bush (Buddleja), Scrophulariaceae Butterwort (Pinguicula), Lentibulariaceae Buxaceae, Buxales Buxales, Eudicots Byblidaceae, Lamiales

Caapi (Banisteriopsis), Malpighiaceae Cabbage (Brassica), Brassicaceae Cabombaceae, Nymphaeales Cactaceae, Caryophyllales Cactus, Cactaceae Caesalpiniaceae, in Fabales Cajuput oil (Melaleuca), Myrtaceae Calabash (Lagenaria), Cucurbitaceae Calabash tree (Crescentia), Bignoniaceae Calamus (Acorus), Acoraceae Calceolaria, Calceolariaceae Calceolariaceae, Lamiales Callicarpa, Lamiaceae Callitrichaceae, in Plantiaginaceae Calochortus, Liliaceae Calophyllaceae, Malpighiales Caltrop (Tribulus), Zygophyllaceae Caltrop, Water (Trapa), Lythraceae Calycanthaceae, Laurales Calyceraceae, Asterales Cambodian mint (Persicaria), Polygonaceae Camellia, Theaceae Camomile (Anthemis, Matricaria, etc), Asteraceae Campanulaceae, Asterales Campanulids, Asterids Camphor/Camphorwood (Cinnamomum),

Lauraceae, and other genera and families Campion (Silene & Lychnis), Caryophyllaceae Campynemataceae, Liliales Canacomyrica, Myricaceae Canadian pondweed (Elodea), Hydrocharitaceae Candleberry (Myrica), Myricaceae Candleberry (Aleurites), Euphorbiaceae Candlenut (Aleurites), Euphorbiaceae Canellaceae, Canellales Canna, Cannaceae Canna, Water (Thalia), Marantaceae Cannabaceae = Cannabidaceae, Rosales Cannabis, Cannabidaceae Cannaceae, Zingiberales Cannonball tree (Couroupita), Lecythidaceae Cape pondweed (Aponogeton), Aponogetonaceae Canotiaceae, in Celastraceae Cape rush (Chondropetalum), Restionaceae Caper (Capparis), Capparaceae Cappar(id)aceae, Brassicales Caprifoliaceae, Dipsacales Capsicum, Solanaceae Carambola (Averrhoa), Oxalidaceae Caraway (Carum), Apiaceae Cardamom (Elattaria & Amomum), Zingiberaceae Cardiopteridaceae, Aquifoliales Cardiopterygaceae = Cardiopteridaceae Cardoon (Cynara), Asteraceae Caricaceae, Brassicales Carlemanniaceae, Lamiales Carnauba (Copernicia), Arecaceae Carnation (Dianthus), Caryophyllaceae Carob (Ceratonia), Fabaceae Carolina Jessamine/Jasmine (Gelsemium),

Gelsemiaceae Carpetweed (Mollugo), Molluginaceae Carpinaceae, in Betulaceae Carrot (Daucus), Apiaceae Caryocaraceae, Malpighiales Cartonemataceae, in Commelinaceae Caryophyllaceae, Caryophyllales Caryophyllales, Eudicots Cascara (Rhamnus), Rhamnaceae Cascarilla (Croton), Euphorbiaceae Cascarilla = Ladenbergia, Rubiaceae Cashew (Anacardium), Anacardiaceae Cassava (Manihot), Euphorbiaceae Cassia (usually Cassia, sometimes Senna or Acacia), Fabaceae Cassia (Cinnamomum), Lauraceae Cassinopsis, Icacinaceae (provisionally) Castor oil (Ricinus), Euphorbiaceae Casuarinaceae, Fagales Catalpa, Bignoniaceae Catnip/Catmint (Nepeta), Lamiaceae Cattail [USA] (Typha), Typhaceae Ceanothus, Rhamnaceae Cecropia, Utricaceae Cecropiaceae, in Urticaceae Cedar, West Indian (Cedrela), Meliaceae Celastraceae, Celastrales Celastrales, Fabids Celery (Apium), Apiaceae Celosia, Amaranthaceae Celtis, Cannabaceae Celtidaceae, in Cannabaceae Centaury (Centaurium), Gentianaceae Centrolepidaceae, Poales Cephalotaceae, Oxalidales Ceratophyllaceae, Ceratophyllales Ceratophyllales, early-branching order Cercidiphyllaceae, Saxifragales Cervantesiaceae, in Santalaceae Ceylon olive (Elaeocarpus), Elaeocarpaceae Champak (Michelia), Magnoliaceae Chaste tree (Vitex), Lamiaceae Chaulmoogra oil (Hydnocarpus), Achariaceae Chenopodiaceae, in Amaranthaceae Cherimoya (Annona), Annonaceae Cherry (Prunus), Rosaceae Cherry, Liberian (Sacoglottis), Humiriaceae Cherry mahogany (Tieghemella), Sapotaceae Chervil (Anthriscus), Apiaceae Chestnut, Horse (Aesculus), Sapindaceae Chestnut, Sweet (Castanea), Fagaceae Chestnut, Water (Eleocharis), Cyperaceae Chia (Salvia), Lamiaceae Chickpea (Cicer), Fabaceae Chicle (Manilkara), Sapotaceae Chicory (Cichorium), Asteraceae Chilli (Capsicum), Solanaceae Chinese tallow tree (Triadica = Sapium),

Euphorbiaceae Chinquapin (Castanea, Chrysolepis &

Castanopsis), Fagaceae Chlaenaceae = Sarcolaenaceae Chloanthaceae, in Lamiaceae Chloranthaceae, Chloranthales Chloranthales, early-branching order Chocolate (Theobroma), Malvaceae Choisya, Rutaceae Chrysanthemum, Asteraceae Chrysobalanaceae, Malpighiales Chumprak (Heritiera), Malvaceae Churnwood (Citronella), Cardiopteridaceae Cicely (Myrrhis), Apiaceae Cineraria, Asteraceae Cinnamon (Cinnamomum), Lauraceae Cinnamon, White (Canella), Canellaceae Cinnamon myrtle (Backhousia), Myrtaceae Ciogue (Nothofagus), Nothofagaceae Circaeasteraceae, Ranunculales Cissus, Vitaceae Cistaceae, Malvales Citronella trees (Citronella), Cardipteridaceae Citrus, Rutaceae Clarkia, Onagraceae Clematis, Ranunculaceae Cleome, Cleomaceae Cleomaceae, Brassicales Clerodendrum, Lamiaceae Clethraceae, Ericales Cleyera, Pentaphylacaceae Clivia, Amaryllidaceae Clover (Trifolium), Fabaceae Cloves (Syzygium), Myrtaceae Clubrush (Scirpus, also Bolboschoenus, Isolepis

& Schoenoplectus), Cyperaceae Clusiaceae, Malpighiales Clutia, Peraceae Cneoraceae, in Rutaceae Cneorum, Rutaceae Coachwood (Ceratopetalum), Cunoniaceae Cobaea, Polemoniaceae Cobaeaceae, in Polemoniaceae Cobra lily (Darlingtonia), Sarraceniaceae Coca, Cocaine (Erythroxylum), Erythroxylaceae Cocoa (Theobroma), Malvaceae Cochlospermaceae, in Bixaceae Cockscomb (Celosia), Amaranthaceae Coco de mer (Lodoicea), Arecaceae Coco-grass (Cyperus), Cyperaceae Coco plum (Chrysobalanus), Chrysobalanaceae Cocoa (Theobroma), Malvaceae Coconut (Cocos), Arecaceae Codonopsis, Campanulaceae Coffee (Coffea), Rubiaceae Coir (Cocos), Arecaceae Cola (Cola), Malvaceae Colchicum, Colchicaceae Colchicaceae, Liliales Coleus (Solenostemon), Lamiaceae Colicwood (Myrsine), Primulaceae Columelliaceae, Bruniales Columnea, Gesneriaceae Comandraceae, in Santalaceae Combretaceae, Myrtales Combretum, Combretaceae Comfrey (Symphytum), Boraginaceae Commelinaceae, Commelinales Commelinales, Monocots (Commelinids) Commelinids, Monocots Congress weed (Parthenium), Asteraceae Connaraceae, Oxalidales Convolvulaceae, Solanales Copal (Copaifera), Fabaceae Corbichonia, Lophiocarpaceae Cordiaceae, in Boraginaceae Cordyline, Asparagaceae Coreopsis, Asteraceae Coriander (Coriandrum), Apiaceae Coriander, Vietnamese (Persicaria), Polygonaceae Coriaria, Coriariaceae Coriariaceae, Cucurbitales Coridaceae, in Primulaceae Cork (Quercus), Fagaceae Corkwood, Florida (Leitneria), Simaroubaceae Corn [USA and elsewhere] (Zea), Poaceae Cornaceae, Cornales Cornales, Asterids Corsiaceae, Liliales Corylaceae, in Betulaceae Corynocarpaceae, Cucurbitales Cosmos, Asteraceae Costa, Costaceae Costaceae, Zingiberales Cotinus, Anacardiaceae Cotton (Gossypium), Malvaceae Coulaceae, Santalales Courgette (Cucurbita), Cucurbitaceae Crape myrtle (Lagerstroemia), Lythraceae Cranberry (Vaccinium), Ericaceae Crane’s bill/Cranesbill (Geranium), Geraniaceae Crassulaceae, Saxifragales Creosote bush (Larrea), Zygophyllaceae Cress, Garden (Lepidium), Brassicaceae Cress, Land (Barbarea), Brassicaceae Cress, Pepper- (Lepidium), Brassicaceae Cress, Water- (Nasturtium), Brassicaceae Cress, Winter (Barbarea), Brassicaceae Crinodendron, Elaeocarpaceae Crocosmia, Iridaceae Crocus, Iridaceae Croomiaceae, in Stemonaceae Crossosomataceae, Crossosomatales Croton, Euphorbiaceae Crowberry (Empetrum), Ericaceae Crypteroniaceae, Myrtales Ctenolophonaceae, Malpighiales Cubeb (Piper), Piperaceae Cucumber (Cucumis), Cucurbitaceae Cucurbitaceae, Cucurbitales Cucurbitales, Fabids Cudjoe wood (Jacquinia = Bonellia), Primulaceae Culver’s root (Veronicastrum), Plantaginaceae Cumin (Cuminum), Apiaceae Cunoniaceae, Oxalidales Cuphea, Lythraceae Curare (Chondrodendron), Menispermaceae Curare (Strychnos), Loganiaceae Currant bush (Ribes), Grossulariaceae Curtisiaceae, Cornales Cuscutaceae, in Convolvulaceae Custard apple (Annona), Annonaceae Cyanastraceae, in Tecophilaeaceae Cyclamen, Primulaceae Cyclanthaceae, Pandanales Cyclocheilaceae, in Orobachaceae Cymodoceaceae, Alismatales Cynomoriaceae – unplaced Cyperaceae, Poales Cypripediaceae, in Orchidaceae Cyphiaceae, in Campanulaceae Cyrillaceae, Ericales Cytinaceae, Malvales

Dabai (Canarium), Burseraceae Daffodil (Narcissus), Amaryllidaceae Dahlia, Asteraceae Daisy (Bellis, Aster, etc), Asteraceae Damiana (Turnera), Passifloraceae Dandelion (Taraxacum), Asteraceae Danta (Nesogordonia), Malvaceae Daphne, Thymelaeaceae Daphniphyllaceae, Saxifragales Dasypogonaceae, Monocots (Commelinids) Date (Phoenix), Arecaceae Date-plum (Diospyros), Ebenaceae Datiscaceae, Cucurbitales Day lily (Hemerocallis), Xanthorrhoeaceae Davidia, Cornaceae Davidiaceae, in Cornaceae Davidsoniaceae, in Cunoniaceae Degame (Calycophyllum), Rubiaceae Degeneriaceae, Magnoliales Desfontainia, Columelliaceae Deutzia, Hydrangeaceae Devil’s claw (Harpagophytum), Pedaliaceae Devil’s claw (Proboscidea), Martyniaceae Dialypetalanthaceae, in Rubiaceae Diapensiaceae, Ericales Diascia, Scrophulariaceae Dicentra, Papaveraceae Dichapetalaceae, Malpighiales Dicrastylidaceae, in Lamiaceae Dictamnus, Rutaceae Didiereaceae, Caryophyllales Didymelaceae, in Buxaceae Diegodendraceae, in Bixaceae Diervillaceae, in Caprifoliaceae Dika nut (Irvingia), Irvingiaceae Dill (Anethum), Apiaceae Dillenia, Dilleniaceae Dilleniaceae, Eudicots Dioncophyllaceae, Caryophyllales Dioscoreaceae, Dioscoreales Dioscoreales, Monocots Dipentodontaceae, Huerteales Dipsacaceae, in Caprifoliaceae Dipsacales, Campanulids Dipterocarpaceae, Malvales Dirachmaceae, Rosales Dissotis, Melastomataceae Ditchgrass (Ruppias), Ruppiaceae Dock (Rumex), Polygonaceae Dodder (Cuscuta), Convolvulaceae Dodecatheon, Primulaceae Dogbane (Apocynum), Apocynaceae Dogwood (Cornus), Cornaceae Donatiaceae, in Styliaceae Doryanthaceae, Asparagales Dove tree (Davidia), Cornaceae Dragon tree (Dracaena), Asparagaceae Dragon’s blood (Croton), Euphorbiaceae Dragon’s blood (Calamus), Arecaceae Dragon’s blood (Daemonorops), Arecaceae Dragon’s blood (Dracaena), Asparagaceae Dragon’s blood (Pterocarpus), Fabaceae Droseraceae, Caryophyllales Drumstick tree (Cassia), Fabaceae Drumstick tree (Moringa), Moringaceae Duckeodendraceae, in Solanaceae Duckweed (Lemna), Araceae Dulongiaceae, in Phyllonomaceae Durian (Durio), Malvaceae Dysphaniaceae, in Amarathaceae

Ebenaceae, Ericales

Ebony (Diospyros), Ebenaceae

Ecdeiocoleaceae, Poales (Echinacea), Asteraceae Eelgrass (Vallisneria), Hydrocharitaceae Eelgrass (Zostera), Zosteraceae Eggfruit (Pouteria), Sapotaceae Ehretiaceae, in Boraginaceae Ekki tree (Lophira), Ochnaceae Elaeagnaceae, Rosales Elaeocarpaceae, Oxalidales Elatinaceae, Malpighiales Elder (Sambucus), Adoxaceae Elder, African forest (Nuxia), Stilbaceae Elderberry (Sambucus), Adoxaceae Elecampane (Inula), Asteraceae Elemi (Canarium), Burseraceae Elephant tree (Bursera), Burseraceae Ellisiophyllaceae, in Plantaginaceae Elm (Ulmus), Ulmaceae Elodea, Hydrocharitaceae Emblingiaceae, Brassicales Emmotum, Icacinaceae (provisionally) Empetraceae, in Ericaceae Emu bush (Eremophilas), Scrophulariaceae Emu bush (Hakea), Proteaceae Endive (Cichorium), Asteraceae Epacridaceae, in Ericaceae Eremolepidaceae, in Santalaceae Eremosynaceae, in Escalloniaceae Ericaceae, Ericales Ericales, Asterids Eriocaulaceae, Poales Erodium, Geraniaceae Erythropalaceae, Santalales Erythroxylaceae, Malpighiales Escallonia, Escalloniaceae Escalloniaceae, Escalloniales Escalloniales, Campanulids Eucalyptus, Myrtaceae Eucommiaceae, Garryales Eucryphia, Cunoniaceae Euphorbiaceae, Malpighiales Euphroniaceae, Malpighiales Eupomatiaceae, Magnoliales Eupteleaceae, Ranunculales Eurya, Pentaphylacaceae Euryalaceae, in Nymphaeaceae Evening primrose (Oenothera), Onagraceae Exacum, Gentianaceae Eyebright (Euphrasia), Orobanchaceae

Fabaceae, Fabales Fabales, Fabids Fabids, Rosids Fagaceae, Fagales Fagales, Fabids Fatsia, Araliaceae Featherfoil (Hottonia), Primulaceae Fennel (Foeniculum), Apiaceae Feverfew (Tanacetum), Asteraceae Fig (Ficus), Moraceae Fig, Hottentot (Carpobrotus), Aizoaceae Fig-marigold (Mesembryanthemum and

other genera), Aizoaceae Figwort (Scrophularia), Scrophulariaceae Firethorn (Pyracantha), Rosaceae Flag, Sweet (Acorus), Acoraceae Flag iris (Iris), Iridaceae Flagellariaceae, Poales Flax (Linum), Linaceae Flax, New Zealand (Phormium), Xanthorrhoeaceae Flindersiaceae, in Rutaceae Flowering rush (Butomus), Butomaceae Flycatcher, Australian (Cephalotus), Cephalotaceae Foetidiaceae, in Lecythidaceae Fogfruit (Phyla), Verbenaceae Forget-me-not (Myosotis), Boraginaceae Forsythia, Oleaceae Fouquieriaceae, Ericales Foxglove (Digitalis), Plantaginaceae Foxglove tree (Paulownia), Paulowniaceae Francoaceae, in Melianthaceae Frangipani (Plumeria), Apocynaceae Frankeniaceae, Caryophyllales Frankincense (Boswellia), Burseraceae Franklin tree (Franklinia), Theaceae Freesia, Iridaceae Fremontodendron, Malvaceae Freijo (Cordia), Boraginaceae Fritillary (Fritillaria), Liliaceae Frogbit (Hydrocharis &

Limnobium), Hydrocharitaceae Fuchsia, Onagraceae Fumariaceae, in Papaveraceae Fumitory (Fumaria), Papaveraceae Fustic, ‘Old’ (Maclura), Moraceae Fustic, ‘Young’ (Cotinus), Anacardiaceae

Gabon nut (Coula), Coulaceae Gaboon (Aucoumea), Burseraceae Gaharu (Aquilaria), Thymelaeaceae Galangal (usually Alpinia, sometimes

Kaempferia, Boesenbergia), Zingiberaceae Galax, Diapensiaceae Galbanum (Ferula), Apiaceae Galingale (Cyperus), Cyperaceae, see also Galangal Galip nut (Canarium), Burseraceae Pili nut (Canarium), Burseraceae Gamboge (Garcinia), Clusiaceae Gardenia, Rubiaceae Garlic (Allium), Amaryllidaceae Garlic-tree (Afrostyrax), Huaceae Garrya, Garryaceae Garryaceae, Garryales Garryales, Lamiids Gaultheria, Ericaceae Geissolomataceae, Celastrales Gentian (Gentiana), Gentianaceae Gentianaceae, Gentianales Gentianales, Lammiids Geosiridaceae, in Iridaceae Geraniaceae, Geraniales Geraniales, Malvids Geranium, Geraniaceae Germander (Teucrium), Lamiaceae Gesneriaceae, Lamiales Gherkin (Cucumis), Cucurbitaceae Gilia, Polemoniaceae Ginger (Zingiber), Zingiberaceae Ginseng (Panax), Araliaceae Gladiolus, Iridaceae Glaucidiaceae, in Ranunculaceae Glaucidium, Ranunculaceae Globe thistle (Echinops), Asteraceae Globularia, Plantaginaceae Globulariaceae, in Plantaginaceae Gloriosa, Colchicaceae Glory bush (Tibouchina), Melastomataceae Gloxinia, Gesneriaceae Goetzeaceae, in Solanaceae Goji berry (Lycium), Solanaceae Golden leaf tree (Chrysophyllum), Sapotaceae Goldenrod (Solidago), Asteraceae Gomortegaceae, Laurales Gonystylaceae, in Thymelaeaceae Goodenia, Goodeniaceae Goodeniaceae, Asterales Gooseberry (Ribes), Grossulariaceae Gooseberry, Cape (Physalis), Solanaceae Gooseberry, Chinese (Actinidia), Actinidiaceae Gooseberry, Indian (Phyllanthus), Phyllanthaceae Gooseberry, Otaheite/Tahitian

(Phyllanthus), Phyllanthaceae Gooseberry, Star (Sauropus), Phyllanthaceae Goosefoots (Chenopodium), Amaranthaceae Gorse (Ulex), Fabaceae Gotu kola (Centella), Apiaceae Goupiaceae, Malpighiales Gourd (various genera inc. Cucumis, Lagenaria, Benincasa, Cucurbita), Cucurbitaceae Granadilla (Passiflora), Passifloraceae Grape (Vitis), Vitaceae Grapefruit (Citrus), Rutaceae Grass (many genera), Poaceae Grass of Parnassus (Parnassia), Celastraceae Grass pinks (Calopogon), Orchidaceae Grass tree (Xanthorrhoea), Xanthorrhoeaceae Grease nut (Hernandia), Hernandiaceae Greengage (Prunus), Rosaceae Greenheart (Chlorocardium), Lauraceae Grevillea, Proteaceae Grey mangrove (Avicenna), Acanthaceae Greyia, Melianthaceae Greyiaceae, in Melianthaceae Griselinia, Griselliniaceae Griseliniaceae, Apiales Gromwell (Lithospermum), Boraginaceae Grossulariaceae, Saxifragales Grubbiaceae, Cornales Guarana (Paullinia), Sapindaceae Guava (Psidium), Myrtaceae Guayule (Parthenium), Asteraceae Guggul gum (Commiphora), Burseraceae Gum trees (Eucalyptus), Myrtaceae Gum guggul (Commiphora), Burseraceae Gunneraceae, Gunnerales Gunnerales, Eudicots Gunpowder-tree (aka Pigeon Wood or Nalita)

(Trema), Cannabaceae Gutta-percha (Palaquium), Sapotaceae Gyrocarpaceae, in Hernandiaceae Gyrostemonaceae, Brassicales

Hackberry (Celtis), Cannabaceae Haemodoraceae, Commelinales Halesiaceae, in Styracaceae Halophytaceae, Caryophyllales Halorag(id)aceae, Saxifragales Hamamelidaceae, Saxifragales Handkerchief tree (Davidia), Cornaceae Hanguanaceae, Commelinales Harmal (Peganum), Nitrariaceae Hawaiian snowbush (Breynia), Phyllanthaceae Hawthorn (Crataegus), Rosaceae Hawthorn, Water (Aponogeton), Aponogetonaceae Hazel (Corylus), Betulaceae Heather/Heath (Calluna, Erica,

Phyllodoce, Cassiope), Ericaceae Heavenly/Sacred bamboo (Nandina), Berberidaceae Hebe (Veronica), Plantaginaceae Hectorellaceae, in Montiaceae Helichrysum, Asteraceae Heliconiaceae, Zingiberales Heliotrope (Heliotropium), Boraginaceae Hellebore (Helleborus), Ranunculaceae Hellebore, False (Adonis), Ranunculaceae Hellebore, False (Veratrum), Melianthaceae Helleborine (Epipactis, & Cephalanthera,

also Calopogon), Orchidaceae Helwingiaceae, Aquifoliales Helxine, Urticaceae Hemerocallidaceae, in Xanthorrhoeaceae Hemlock (Conium), Apiaceae Hemlock, Water (Cicuta), Apiaceae Hemlock water dropwort (Oenanthe), Apiaceae Hemp (Cannabis), Cannabidaceae Hemp, African (Sparrmannia), Malvaceae Hemp, Manila (Musa), Musaceae Henbane (Hyoscyamus), Solanaceae Henna (Lawsonia), Lythraceae Henriqueziaceae, in Rubiaceae Hernandiaceae, Laurales Heteropyxidaceae, in Myrtaceae Heuchera, Saxifragaceae Hibbertia, Dilleniaceae Hibiscus, Malvaceae Hickory (Carya), Juglandaceae Hickory, Beaked (Annamocarya), Juglandaceae Himantandraceae, Magnoliales Hippocastanaceae, in Sapindaceae Hippocrateaceae, in Celastraceae Hippuridaceae, in Plantaginaceae Hog plum (Ximenia), Ximeniaceae Hog plum (Spondias), Anacardiaceae Holly (Ilex), Aquifoliaceae Holly, Sea (Acanthus), Acanthaceae Holly, Sea (Eryngium), Apiaceae Hollyhock (Alcea), Malvaceae Honesty (Lunaria), Brassicaceae Honey berry (Lonicera), Caprifoliaceae Honey flower (Lambertia), Proteaceae Honey flower (Melianthus), Melianthaceae Honey myrtle (Melaleuca), Myrtaceae Honeysuckle (Lonicera), Caprifoliaceae Hop (Humulus), Cannabidaceae Hop-hornbeam (Ostrya), Betulaceae Hoplestigmataceae, in Boraginaceae (provisionally) Hornbeam (Carpinua), Betulaceae Hornwort (Ceratophyllum), Ceratophyllaceae Horse chestnut (Aesculus), Sapindaceae Hortoniaceae, in Monimiaceae Hosta, Asparagaceae Hottentot fig (Carpobrotus), Aizoaceae Houmiriaceae = Humiriaceae House-leek (Sempervivum), Crassulaceae Hoya, Apocynaceae Hua, Huaceae Huaceae, Fabids Hualo (Nothofagus), Nothofagaceae Huambula wood (Minquartia), Coulaceae Huckleberry (Vaccinium, Gaylussacia), Ericaceae Huckleberyy, Garden (Solanum), Solanaceae Hugoniaceae, in Linaceae Humiriaceae, Malpighiales Hyacinth (Hyacinthus), Asparagaceae Hyacinth, Water (Eichhornia), Pontederiaceae Hyacinthaceae, in Asparagaceae Hydatellaceae, Nympaeales Hydnoraceae, Piperales Hydrangea, Hydrangeaceae Hydrangeaceae, Cornales Hydrastis, Ranunculaceae Hydrocharitaceae, Alismatales Hydrocotylaceae, in Araliaceae Hydrophyllaceae, in Boraginaceae Hydrostachyaceae, Cornales (provisionally) Hypecoum, Papaveraceae Hypericaceae, Malpighiales Hypericum, Clusiaceae Hypoxidaceae, Asparagales Hypseocharis, Geraniaceae Hyssop (Hyssopus & Agastache), Lamiaceae

Icacinaceae, Lamiids

Idigbo (Terminalia), Combretaceae

Illiciaceae, in Schisandraceae Imbuia (Phoebe), Lauraceae Impatiens, Balsamaceae Indian shot (Canna), Cannaceae Indian mast tree (Polyalthia), Annonaceae Indian pipe (Monotropa), Ericaceae Indigo (Indigofera), Fabaceae Inoy/Inoi nut Poga, Anisophylleaceae Ipecacuanha (Psychotria), Rubiaceae Ipecacuanha, White (Ionidium = Hybanthus),

Violaceae Iridaceae, Asparagales Iris, Iridaceae Iroko (Milicia or Chlorophora = Maclura), Moraceae Ironwood - many genera and families Irvingiaceae, Malpighiales Isano oil (Ongokea), Aptandraceae Ivy (Hedera), Araliaceae Ivy, Poison (Toxicodendron = Rhus), Anacardiaceae Ixioliriaceae, Asparagales Ixonanthaceae, Malpighiales

Jaborandi (Pilocarpus), Rutaceae Jaboty (Erisma), Vochysiaceae Jaboty (Helicostylis), Moraceae Jacaranda, Bignoniaceae Jackfruit (Artocarpus), Moraceae Jacob’s ladder (Polemonium), Polemoniaceae Jalop (Ipomoea), Convolvulaceae Jasmine (Jasminum), Oleaceae Jasmine, Cape (Gardenia), Rubiaceae Jasmine, New Zealand (Parsonsia), Apocynaceae Jasmine, Night-blooming (Cestrum), Solanaceae Jasmine, Night-flowering (Nyctanthes), Oleaceae Jasmine, Star (Trachelospermum), Apocynaceae Jasmine tree (Millingtonia), Bignoniaceae Jatropha, Euphorbiaceae Jelutong (Dyera), Apocynaceae Jessamine/Jasmine, Carolina (Gelsemium),

Gelsemiaceae Jimson weed (Datura), Solanaceae Jobo (Spondias), Anacardiaceae Joinvilleaceae, Poales Jojoba (Simmondsia), Simmondsiaceae Jongkong (Dactylocladus), Crypteroniaceae Joshua tree (Yucca), Asparagaceae Jovellana, Calceolariaceae Judas tree (Cercis), Fabaceae Juglandaceae, Fagales Jujube (Ziziphus), Rhamnaceae Julianiaceae, in Anacardiaceae Juncaceae, Poales Juncaginaceae, Alismatales Jute (Corchorus), Malvaceae

Kadsura, Schisandraceae Kalanchoe, Crassulaceae Kalmia, Ericaceae Kamahi (Weinmannia), Cunoniaceae Kangaroo paw (Anigozanthus & Macropidia),

Haemodoraceae Kapok (Ceiba), Malvaceae Kapur (Dryobalanops), Dipterocarpaceae Karanda (Carissa), Apocynaceae Karanda nut (Elaeocarpus), Elaeocarpaceae Karir (Capparis), Capparaceae Karo tree (Pittosporum), Pittosporaceae Katsura tree (Cercidiphyllum), Cercidiphyllaceae Kapok (Ceiba), Malvaceae Kava (Piper), Piperaceae Keaki (Zelkova), Ulmaceae Kei apple (Dovyalis = Aberia), Salicaceae Khat (Catha), Celastraceae Kirkiaceae, Sapindales Kite tree (Nuxia), Stilbaceae Kiwi fruit (Actinidia), Actinidiaceae Knapweed (Centuarea), Asteraceae Kohekohe (Dysoxylum), Meliaceae Koeberliniaceae, Brassicales Kokoon (Kokoona), Celastraceae Kola nut, see Cola Krabak (Anisoptera), Dipterocarpaceae Krameriaceae, Zygophyllales Kudzu (Pueraria), Fabaceae Kulim wood (Scorocarpus), Strombosiaceae Kumquat (Fortunella), Rutaceae

Labdanum (Cistus), Cistaceae Labrador tea (Ledum), Ericaceae Laburnum, Fabaceae Lacistemataceae, Malpigiales Lacquer (Rhus), Anacardiaceae Lactoridaceae, Piperales Lamiaceae, Lamiales Lamiales, Lamiids Lamiids, Asterids Lanariaceae, Asparagales Lantana, Verbenaceae Lapageria, Philesiaceae Lardizabalaceae, Ranunculales Lasora (Cordia), Boraginaceae Lauraceae, Laurales Laurales, Magnoliids Laurel, Azores (Laurus), Lauraceae Laurel, Bay (Laurus), Lauraceae Laurel, Californian (Umbellularia), Lauraceae Laurel, Camphor (Cinnamomum), Lauraceae Laurel, Cherry (Prunus), Rosaceae Laurel, Great (Rhododendron), Ericaceae Laurel, Mountain (Kalmia), Ericaceae Laurel, New Zealand (shrub) (Coprosma),

Rubiaceae Laurel, New Zealand (tree) (Corynocarpus),

Corynocarpaceae Laurel, Portugal (Prunus), Rosaceae Laurel, Spotted (Aucuba), Garryaceae Laurel, Sumac (Malosma = Rhus), Anacardiaceae Laurel, Tasmanian (Anopterus), Escalloniaceae Lavender (Lavandula), Lamiaceae Lavender, Sea (Limonium), Plumbaginaceae Laxmanniaceae, in Asparagaceae Leadwood (Krugiodendron), Rhamnaceae Leadwood (Ceratostigma), Plumbaginaceae Leadwood tree (Combretum), Combretaceae Leadwort (Ceratostigma), Plumbaginaceae Leafflower (Phyllanthus), Phyllanthaceae Leatherwood (Cyrilla), Cyrillaceae Leatherwood (Dirca), Thymelaeaceae Leatherwood (Eucryphia), Cunoniaceae Leatherwood (Olearia), Asteraceae Lebombo ironwood (Androstachys), Picrodendraceae Lechenaultia = Leschenaultia Lecythidaceae, Ericales Ledocarpaceae, in Vivianiaceae Leeaceae, in Vitaceae Leek (Allium), Amaryllidaceae Leitneriaceae, in Simaroubaceae Lemnaceae, in Araceae Lemon (Citrus), Rutaceae Lemon thorn (Cassinopsis), Icacinaceae (provisionally) Lenga (Nothofagus), Nothofagaceae Lennoaceae, in Boraginaceae Lentibulariaceae, Lamiales Lentil (Lens), Fabaceae Leptaulus, Cardiopteridaceae Leschenaultia, Goodeniaceae Lettuce (Lactuca), Asteraceae Lettuce, Sea (Scaevola), Goodeniaceae Lewisia, Montiaceae Leycesteria, Caprifoliaceae Lichee, see Lychee Lignum vitae (Guaiacum), Zygophyllaceae Lilac (Syringa), Oleaceae Lilaeaceae, in Juncaginaceae Liliaceae, Liliales Liliales, Monocots Lily, Liliaceae Lily, Peruvian (Alstroemera), Alstroemeriaceae Lily-of-the -valley (Convallaria), Asparagaceae Lily-of-the-valley tree (Clethra), Clethraceae Lily-of-the-valley tree (Crinodendron), Elaeocarpaceae Lily-of-the-valley tree (Oxydendron), Ericaceae Lime (Citrus), Rutaceae Lime tree [UK] (Tilia), Malvaceae Limnanthaceae, Brassicales Limnocharitaceae, in Alismataceae Linaceae, Malpighiales Linaloa, Burseraceae Linderniaceae Linnaeaceae, in Caprifoliaceae Linseed (Linum), Linaceae Liquidambar, Altingiaceae Liquorice/Licorice (Glycyrrhiza), Fabaceae Lisianthus (Eustoma), Gentianaceae Lissocarpaceae, in Ebenaceae Litchi, see Lychee Lithops, Aizoaceae Littorella, Plantaginaceae Lizard’s tail (Saururus), Saururaceae Loasaceae, Cornales Lobelia, Campanulaceae Lobeliaceae, in Campanulaceae Loblolly bay (Gordonia), Theaceae Lobster-claw (Clianthus), Fabaceae Lobster-claw (Heliconia), Heliconiaceae Lodh (bark) (Symplocos), Symplocaceae Loganiaceae, Gentianales Longan (Dimocarpus), Sapindaceae Loosestrife, purple (Lythrum), Lythraceae Loosestrife, yellow (Lysimachia), Primulaceae Lophopyxidaceae, Malpighiales Lopseed (Phryma), Phrymaceae Loquat (Eriobotrya), Rosaceae Loranthaceae, Santalales Lotus, Sacred (Nelumbo), Nelumbaceae

(also Nymphaea, Nymphaeaceae) Lousewort (Pedicularis), Orobanchaceae Lovage (Levisticum), Apiaceae Lovegrass (Eragrostis), Poaceae Lowiaceae, Zingiberales Luffa, Cucurbitaceae Lungwort (Pulmonaria), Boraginaceae Lupin (Lupinus), Fabaceae Luzuriagaceae, Liliales Lychee/Lichee/Litchi (Litchi), Sapindaceae Lythraceae, Myrtales Lythrum, Lythraceae

Macachi (Arjona), Schoepfiaceae Macadamia, Proteaceae Macary bitter (Picramnia), Picramniaceae Mace (Myristica), Myristicaceae Mackinlayaceae, in Apiaceae Madagascar periwinkle (Catharanthus),

Apocynaceae Madder (Rubia), Rubiaceae Madeira vine (Anredera), Basellaceae Madrone (Arbutus), Ericaceae Madwort (Asperugo), Boraginaceae Madwort (Lobularia, Aurinia and Alyssum), Brassicaceae Maesaceae, in Primulaceae Magnolia, Magnoliaceae Magnoliaceae, Magnoliales Magnoliales, Magnoliids Maguey (Agave), Asparagaceae Mahogany (Swietenia), True m., West

Indian m., Cuban m. & Honduras m., Meliaceae, (also species in Khaya & Entandrophragma in

Meliaceae)

Mahogany, African (Khaya), Meliaceae Mahogany, Chinese (Toona), Meliaceae Mahogany, New Zealand (Dysoxylum), Meliaceae Maize (Zea), Poaceae Makita (Atuna), Chrysobalanaceae Makore (Tieghemella), Sapotaceae Malesherbiaceae, in Passifloraceae Mallow (Malva), Malvaceae Malpighiaceae, Malpighiales Malpighiales, Fabids Malvaceae, Malvales Malvales, Malvids Malvids, Rosids Mamaki (Pipturus), Utricaceae Mamey sapote (Pouteria), Sapotaceae Mamoncilla (Melicoccus), Sapindaceae Manatee grass (Syringodium & Cymodocea),

Cymodoceaceae Mandrake (Mandragora), Solanaceae Mangeao (Litsea), Lauraceae Mango (Mangifera), Anacardiaceae Mango, Bush (Irvingia), Irvingiaceae Mangosteen (Garcinia), Clusiaceae Mangrove, Black (Avicenna), Acanthaceae Mangrove, Red (Rhizophora), Rhizophoraceae Mangrove, White/Grey (Avicenna), Acanthaceae Mangrove, White (Conocarpus, Languncularia &

Lumnitzera), Combretaceae Mangrove apple (Sonneratia), Lythraceae Mangrove palm (Nypha), Arecaceae Manila hemp (Musa), Musaceae Manwood, Black (Minquartia), Coulaceae Maple (Acer), Sapindaceae Maple, Queensland (Flindersia), Rutaceae Maracaibo boxwood (Casearia), Salicaceae Marantaceae, Zingiberales Marcgraviaceae, Ericales Mare’s tail (Hippuris), Plantaginaceae Marigold (Tagetes), Asteraceae Marigold, Corn (Glebionis), Asteraceae Marigold, Desert (Baileya), Asteraceae Marigold, Marsh (Caltha), Ranunculaceae Marigold, Mexican (Tagetes), Asteraceae Marigold, Pot (Calendula), Asteraceae Marigold, Tree (Tithonia), Asteraceae Marjoram (Origanum), Lamiaceae Marrow (Cucurbita), Cucurbitaceae Marshmallow (Althaea), Malvaceae Martyniaceae, Lamiales Mast tree, Indian (Polyalthia), Annonaceae Mastic (Pistacia), Anacardiaceae Maundia, Juncaginaceae Mayacaceae, Poales Mayten (Maytenus), Celastraceae Mbura (Parinari), Chrysobalanaceae Meadow saffron (Colchicum), Colchicaceae Meadowfoam (Limnanthes), Limnanthaceae Medinilla, Melastomataceae Medusagyne, Ochnaceae Medusandraceae, in Peridiscaceae Melanophyllaceae, in Torricelliaceae Melanthiaceae, Liliales Melastomataceae, Myrtales Meliaceae, Sapindales Melianthaceae, Geraniales Meliosmaceae, in Sabiaceae Melon (Cucumis), Cucurbitaceae Melon, Bitter (Momordica), Cucurbitaceae Melon, Horned (Cucumis), Cucurbitaceae Melon, Water (Citrullus), Cucurbitaceae Melon, Winter (Benincasa), Cucurbitaceae Mendonciaceae, in Acanthaceae Mengkulang (Heritiera), Malvaceae Menispermaceae, Ranunculales Menyanthaceae, Asterales Meranti (Shorea), Dipterocarpaceae Mercury (Mercurialis), Euphorbiaceae Mersawa (Anisoptera), Dipterocarpaceae Mescal (Agave), Asparagaceae Mesembryanthemum, Aizoaceae Mesquite (Prosopis), Fabaceae Metteniusaceae, Lamiids Micranthemum, Linderniaceae Mignonette (Reseda), Resedaceae Milk-vetch (Astragalus), Fabaceae Milkweed (Asclepias), Apocynaceae Milkweed (Sonchus), Asteraceae Milkwort (Polygala), Polygalaceae Millet (various genera), Poaceae Mimosa, Fabaceae Mint, including Spearmint, Peppermint &

Horse mint (Mentha), Lamiaceae Mint, Cambodian (Persicaria), Polygonaceae Mint bush (Prostanthera), Lamiaceae Mirabilis, Nyctaginaceae Misodendraceae, Santalales Mistletoe, various genera and families in Santalales,

including European m. (Viscum), American m. (Phoradendron) and Dwarf m. (Arceuthobium),

all in Viscaceae Mitrastemonaceae = Mitrastem(m)ataceae,

Ericales (provisionally) Mock orange, (Philadelphus) Hydrangeaceae Molluginaceae, Caryophyllales Mombin (Spondias), Anacardiaceae Monimiaceae, Laurales Monkey flower (Mimulus), Phrymaceae Monotropa, Ericaceae Monotropaceae, in Ericaceae Monstera, Araceae Moonseed (Cocculus & Menispermum),

Menispermaceae Moraceae, Rosales Morinaceae, in Caprifoliaceae Moringaceae, Brassicales Morning glory (Ipomoea, and other genera),

Convolvulaceae Moschatel (Adoxa), Adoxaceae Moss-rose (Portulaca), Portulacaceae Moutan (Paeonia), Paeoniaceae Mpingo wood (Dalbergia), Fabaceae Muhuhu oil/wood (Brachyleana), Asteraceae Mulberry (Morus), Moraceae Mullein (Verbascum), Scrophulariaceae Mulukhiyah (Corchorus), Malvaceae Muntingiaceae, Malvales Musaceae, Zingiberales Mustard (Sinapis & Brassica), Brassicaceae Mustard-tree (Nicotiana), Solanaceae Mustard-tree (Salvadora), Salvadoraceae Myodocarpaceae, Apiales Myoporaceae, in Scrophulariaceae Myricaceae, Fagales Myristicaceae, Magnoliales Myrothamnaceae, Gunnerales Myrrh (Commiphora), Burseraceae Myrrh, Sweet (Opopanax), Apiaceae Myrsinaceae, in Primulaceae Myrtaceae, Myrtales Myrtales, Malvids Myrtle (Myrtus), Myrtaceae Myrtle, Aniseed (Syzygium), Myrtaceae Myrtle, Beech (Nothofagus), Nothofagaceae Myrtle, Black (Diospyros), Ebenaceae Myrtle, Brown (Choricarpia), Myrtaceae Myrtle, Cinnamon (Backhousia), Myrtaceae Myrtle, Crape (Lagerstroemia), Lythraceae Myrtle, Creeping (Vinca), Apocynaceae Myrtle, Dwarf Australian (Myoporum),

Scrophulariaceae Myrtle, Honey (Melaleuca), Myrtaceae Myrtle, Lemon (Backhousia), Myrtaceae Myrtle, Rose (Archirhodomyrtus), Myrtaceae Myrtle, Wax (Myrica), Myricaceae Myrtle, White (Auranticarpa), Pittosporaceae Myrtle, White (Hypocalymma), Myrtaceae Myrtle, Willow (Agonis), Myrtaceae Myrtle hakea (Hakea), Proteaceae Myrtle wattle (Acacia), Fabaceae

Najadaceae, in Hydrocharitaceae Nanodeaceae, in Santalaceae Nartheciaceae, Dioscoreaceae Nasturtium, Garden (Tropaeolum), Tropaeolaceae Natal bottlebrush (Greyia), Melianthaceae Nectarine (Prunus), Rosaceae Neem (Azadirachta), Meliaceae Nelumbonaceae, Proteales Nemesia, Scrophulariaceae Nemophila, Boraginaceae Nepenthaceae, Caryophyllales Nesogenaceae, in Orobanchaceae Nettle (Urtica), Urticaceae Nettle, Rock (Eucnide), Loasaceae Neuradaceae, Malvales Ngaio (Myoporum), Scrophulariaceae Nicobar breadfruit (Pandanus), Pandanaceae Nightshade, Deadly (Atropa), Solanaceae Nightshade, Stinking (Hyoscyamus), Solanaceae Nightshade, Woody (Solanum), Solanaceae Nire (Nothofagus), Nothofagaceae Nitrariaceae, Sapindales Nolanaceae, in Solanaceae Nothofagaceae, Fagales Nutmeg (Myristica), Myristicaceae Nyatah (Palaquium), Sapotaceae Nyctaginaceae, Caryophyllales Nymphaeaceae, Nymphaeales Nymphaeales, early-branching order Nymphoides, Menyanthaceae Nyssaceae, in Cornaceae

Oak (Quercus), Fagaceae Oak, Tan (Lithocarpus), Fagaceae Oat (Avena), Poaceae Ombu (Phytolacca), Phytolaccaceae Oca (Oxalis), Oxalidaceae Oca quina (Ullucus), Basellaceae Ochna, Ochnaceae Ochnaceae, Malpighiales Ocotillo (Fouquieria), Fouquieriaceae Octoknemaceae, Santalales Okra (Abelmoschus), Malvaceae Olacaceae, Santalales Oleaceae, Lamiales Oleander (Nerium), Apocynaceae Olearia, Asteraceae Oleaster (Elaeagnus), Elaeagnaceae Oliniaceae, in Penaeaceae Olive (Olea), Oleaceae Olive, Ceylon (Elaeocarpus), Elaeocarpaceae Olivillo (Aextoxicon), Aextoxicaceae Ombu (Phytolacca), Phytolaccaceae Omu (Entandrophragma), Meliaceae Onagraceae, Myrtales Oncothecaceae, Lamiids Onion (Allium), Amaryllidaceae Opepe (Nauclea), Rubiaceae Opiliaceae, Santalales Opium (Papaver), Papaveraceae Opopanax, Apiaceae Orache (Atriplex), Amaranthaceae Orange (Citrus), Rutaceae Orchid (many genera), Orchidaceae Orchidaceae, Asparagales Oregano (Origanum), Lamiaceae Oregano, Mexican (Lippia), Verbenaceae Orobanchaceae, Lamiales Osage orange (Maclura), Moraceae Osier (Salix), Salicaceae Osmanthus, Oleaceae Otaheite gooseberry (Phyllanthus), Phyllanthaceae Ovoga wood (Poga), Anisophylleaceae Oxalidaceae, Oxalidales Oxalidales, Fabids

Pachysandra, Buxaceae Padauk wood (Pterocarpus), Fabaceae Paeoniaceae, Saxifragales Pagoda tree (Styphnolobium), Fabaceae Palm (many genera), Arecaceae Palm oil (Elaeis), Arecaceae Palmyra palm (Borassus), Arecaceae Panama hat plant (Carludovica), Cyclanthaceae Pandaceae, Euphorbiales Pandanaceae, Malpighiales Pansy (Viola), Violaceae Papaveraceae, Ranunculales Papaw (Asimina), Annonaceae Papaya (Carica), Caricaceae Papyrus (Cyperus), Cyperaceae Paracryphiaceae, Paracryphiales, Paracryphiales, Campanulids Paradise nut (Lecythis), Lecythidaceae Parchment-bark (Pittosporum), Pittosporaceae Paris, Melanthiaceae Parnassia, Celastraceae Paropsia, Passifloraceae Parsley (Petroselinum), Apiaceae Parsnip (Pastinaca), Apiaceae Passifloraceae, Malpighiales Passion flower (Passiflora), Passifloraceae Passionfruit (Passiflora), Passifloraceae Patchouli (Pogostemon), Lamiaceae Pau marfin (Balfourodendron), Rutaceae Paulownia, Paulowniaceae Paulowniaceae, Lamiales Pawpaw (Asimina), Annonaceae Pea (Pisum), Fabaceae Pea, sweet (Lathyrus), Fabaceae Peach (Prunus), Rosaceae Peanut (Archis), Fabaceae Pear (Pyrus), Rosaceae Pear, African (Dacryodes), Burseraceae Pear, African (Manilkara), Sapotaceae Pear, Prickly (Opuntia), Cactaceae Pearwood, African (Baillonella), Sapotaceae Pearlwort (Sagina & Colobanthus), Caryophyllaceae Pecan (Carya), Juglandaceae Pedaliaceae, Lamiales Peepul (Ficus), Moraceae Pelargonium, Geraniaceae Pellicieraceae, in Tetrameristaceae Pellitory-of-the-wall (Parietaria), Urticaceae Peltanthera, Gesneriaceae Penaeaceae, Myrtales Pennantiaceae, Apiales Pennyroyal (Mentha), Lamiaceae Pennywort, Water (Hydrocotyle), Araliaceae Pentadiplandraceae, Brassicales Pentaphragmataceae, Asterales Pentaphylacaceae, Ericales Peony (Paeonia), Paeoniaceae Peperomiaceae, in Piperaceae Pepper (Piper), Piperaceae Pepper (Capsicum), Solanaceae Pepper, Betel (Piper), Piperaceae Pepper, Sichuan/Szechuan (Zanthoxylum),

Rutaceae Pepper, Tasmanian (Drimys = Tasmannia),

Winteraceae Pepper elder (Piper & Peperomia), Piperaceae Peppercress (Lepidium), Brassicaceae Pequi (Caryocar), Caryocaraceae Pera, Peraceae Peraceae, Malpighiales Peroba, White (Paratecoma), Bignoniaceae Petai (Parkia), Fabaceae Peridiscaceae, Saxifragales Perilla, Lamiaceae Periwinkle (Vinca), Apocynaceae Periwinkle, Madagascar (Catharanthus),

Apocynaceae Petermanniaceae, Liliales Peroba rosa (Aspidosperma) , Apocynaceae Persimmon (Diospyros), Ebenaceae Petiveria, Phytolaccaceae Petroleum nut (Pittosporum), Pittosporaceae Petrosaviaceae, Monocots Petunia, Solanaceae Peyote (Lophophora), Cactaceae Phacelia, Boraginaceae Phellinaceae, Asterales Philadelphus, Hydrangeaceae Philesiaceae, Liliales Philodendron, Araceae Philydraceae, Commelinales Phlox, Polemoniaceae Phormium, Xanthorrhoeaceae Phrym(at)aceae, Lamiales Phyllanthaceae, Malpighiales Phyllonomaceae, Aquifoliales Physenaceae, Caryophyllales Phytolaccaceae, Caryophyllales Pickrel weed (Pontederia), Pontederiaceae Picramniaceae, Picramniales Picramniales, Malvids Picrodendraceae, Malpighiales Pieris, Ericaceae Pilea, Urticaceae Pili nut (Canarium), Burseraceae Pimpernel (Anagallis), Primulaceae Pineapple (Ananas), Bromeliaceae Pinedrops (Pterospora), Ericaceae Pink (Dianthus), Caryophyllaceae Piperaceae, Piperales Piperales, Magnoliids Pipevine (Aristolochia), Aristolochiaceae Pipewort (Eriocaulon), Eriocaulaceae Pistachio (Pistacia), Anacardiaceae Pitchapple (Clusia), Clusiaceae Pitcher plants, several families: Nepenthaceae,

Sarraceniaceae & Cephalotaceae Pittosporum, Pittosporaceae Pittosporaceae, Apiales Plane trees (Platanus), Platanaceae Plantaginaceae, Lamiales Plantain (Musa), Musaceae Plantain (Plantago), Plantaginaceae Plantain, Floating water (Luronium), Alismataceae Plantain, Water (Alisma), Alismataceae Plantago, Plantaginaceae Platanaceae, Proteales Platycarpum, Rubiaceae Platycodon, Campanulaceae Plocospermataceae, Lamiales Plum (Prunus), Rosaceae Plum, Australian native, (Sideroxylon ), Sapotaceae Plum, Yellow (Ximenia), Ximeniaceae Plumbaginaceae, Caryophyllales Plumbago, Plumbaginaceae Poaceae, Poales Poached egg plant (Limnanthes), Limnanthaceae Pod grass (Scheuchzeria), Scheuchzeriaceae Podophyllaceae, in Berberidaceae Podostemaceae, Malpighiales Poison ivy (Toxicodendron = Rhus), Anacardiaceae Poisonleaf (Dichapetalum), Dichapetalaceae Pokeweed (Phytolacca), Phytolaccaceae Polemoniaceae, Ericales Polemonium, Polemoniaceae Polygalaceae, Fabales Polygonaceae, Polygonales Polyosomaceae, in Escalloniaceae Pomegranate (Punica), Lythraceae Pomelo (Citrus), Rutaceae Pondweed (Potamogeton), Potamogetonaceae Pondweed, Canadian (Elodea), Hydrocharitaceae Pondweed, Cape (Aponogeton), Aponogetonaceae Pontederiaceae, Commelinales Popcorn-flower (Cryptantha & Plagiobothrys),

Boraginaceae Poplar (Populus), Salicaceae Poplar, Balsam (Populus), Salicaceae Poppy (Papaver & Meconopsis), Papaveraceae Poppy, California (Eschscholzia), Papaveraceae Poppy, Desert bearpaw (Arctomecon),

Papaveraceae Poppy, Horned (Glaucium), Papaveraceae Poppy, Matilija (Romneya), Papaveraceae Poppy, Nepal (Meconopsis), Papaveraceae Poppy, Pygmy (Canbya), Papaveraceae Poppy, Prickly (Argemone), Papaveraceae Poppy, Tree (Dendromecon & Romneya), Papaveraceae Poppy, Tulip (Hunnemannia), Papaveraceae Poppy, Water (Hydrocleys), Alismataceae Poppy, Welsh (Meconopsis), Papaveraceae Poppy, Wind (Stylomecon), Papaveraceae Portulacaceae, Caryophyllales Posidoniaceae, Alismatales Potaliaceae, in Gentianaceae Potamogetonaceae, Alismatales Potato (Solanum), Solanaceae Potato, Sweet (Ipomoea), Convolvulaceae Pottingeriaceae, in Celastraceae Prayer plant (Maranta & Calathea), Marantaceae Prickly pear (Opuntia), Cactaceae Primrose (Primula), Primulaceae Primulaceae, Ericales Privet (Ligustrum), Oleaceae Protea, Proteaceae Proteaceae, Proteales Proteales, Early-branching Eudicot Psiloxylaceae, in Myrtaceae Ptaeroxylaceae, in Rutaceae Pulque (Agave), Asparagaceae Pumpkin (Cucurbita), Cucurbitaceae Punicaceae, in Lythraceae Punnai oil (Calophyllum), Calophyllaceae Purpleheart (Peltogyne) , Fabaceae Purslane, Summer (Portulaca), Portulaceae Purslane, Winter (Claytonia), Montiaceae Purslane, Pink (Claytonia), Montiaceae Purslane, Sea (Halimione), Amaranthaceae Putranjivaceae, Malpighiales Pyrolaceae, in Ericaceae Quaruba (Vochysia), Vochysiaceae Quassia, Simaroubaceae Queensland ‘walnut’ (Endiandra), Lauraceae Quiinaceae, in Ochnaceae Quillajaceae, Fabales Quince (Cydonia & Pseudocydonia), Rosaceae Quince, Bengal (Aegle), Rutaceae Quince, Flowering (Chaenomeles), Rosaceae Quinine (Cinchona), Rubiaceae Quinoa (Chenopodium), Amaranthaceae Quintinia, Paracryphiaceae

Radish (Raphanus), Brassicaceae

Raffia (Raphia), Arecaceae Rafflesia, Rafflesiaceae Rafflesiaceae, Malpighiales Ragwort (Senecio), Asteraceae Rain tree (Albizia), Fabaceae Rambutan (Nephelium), Sapindaceae Ramie fibre (Boehmeria), Urticaceae Ramin (Gonystylus), Thymelaeaceae Rampion (Phyteuma & Campanula), Campanulaceae Ranunculaceae, Ranunculales Ranunculales, Early-branching Eudicot Raoulia, Asteraceae Rapateaceae, Poales Rape/Rapeseed (Brassica), Brassicaceae Rasamala (Altingia), Altingiaceae Rattan (various genera), Arecaceae Rattle (Rhinanthus), Orobanchaceae Rauli (Nothofagus), Nothofagaceae Red mangrove (Rhizophora), Rhizophoraceae Red-hot poker (Kniphofia), Xanthorrhoeaceae Reed (Phragmites, Arundo, etc.), Poaceae Reedmace (Typha), Typhaceae Rehmannia, Orobanchaceae (provisionally) Resedaceae, Brassicales Restionaceae, Poales Retziaceae, in Stilbaceae Rhabdodendraceae, Caryophyllales Rhamnaceae, Rosales Rhatany root (Krameria), Krameriaceae Rhipogonaceae, Liliales Rhizophoraceae, Malpighiales Rhododendron, Ericaceae Rhoeo = Tradescantia, Commelinaceae Rhoipteleaceae, in Juglandaceae Rhopalocarpaceae, in Sphaerosepalaceae Rhubarb (Rheum), Polygonaceae Rhynchocalycaceae, in Penaeaceae Rhynchotheca, Vivianiaceae Rice (Oryza), Poaceae River oak (Casuarina), Casuarinaceae Riverweed (Podostemon), Podostemaceae Roble (Nothofagus), Nothofagaceae Rock nettle (Eucnide), Loasaceae Rockrose (Cistus et al.), Cistaceae Rodgersia, Saxifragaceae Roridulaceae, Ericales Rosaceae, Rosales Rosales, Fabids Rose (Rosa), Rosaceae Rose imperial (Cochlospermum), Bixaceae Rosemary (Rosmarinus), Lamiaceae Rosewood, various genera, mainly Dalbergia, Fabaceae

including Brazilian rosewood & (East) Indian rosewood

Rosewood, Burmese (Pterocarpus), Fabaceae Rosids, Eudicots Rousseaceae, Asterales Rubber (Hevea), Euphorbiaceae Rubber tree (Ficus), Moraceae Rubber tree, Panama (Castilla), Moraceae Rubiaceae, Gentianales Rudbeckia, Asteraceae Rue (Ruta), Rutaceae Ruppiaceae, Alismatales Ruscus, Asparagaceae Ruscaceae, in Asparagaceae Rush (Juncus), Juncaceae Rush, Cape (Chondropetalum), Restionaceae Rush, Flowering (Butomus), Butomaceae Russian vine (Fallopia), Polynogaceae Rutaceae, Sapindales Rye (Secale), Poaceae

Sabiaceae, Early-branching Eudicots Saccifoliaceae, in Gentianaceae Sacred bamboo (Nandina), Berberidaceae Sacred lotus, see Lotus Saffron (Crocus), Iridaceae Saffron, Meadow (Colchicum), Colchicaceae Sage (Salvia), Lamiaceae Sage, Antelope (Eriogonum), Polygonaceae Sage, Jerusalem (Phlomis), Lamiaceae Sage, Russian (Perovskia), Lamiaceae Sage, Texas (Leucophyllum), Plantaginaceae Sage, Wood (Teucrium), Lamiaceae Sagebrush (Artemisia), Asteraceae Sago (Metroxylon), Arecaceae Saguaro (Carnegiea), Cactaceae St. John’s wort (Hypericum), Clusiaceae Sakaki (Cleyera), Pentaphylacaceae Salep (Orchis), Orchidaceae Salicaceae, Malpighiales Sallow (Salix), Salicaceae Saltbush (Atriplex), Amaranthaceae Saltwort (Batis), Bataceae Saltwort (Salicornia & Salsola), Amaranthaceae Salvadoraceae, Brassicales Samphire, Golden (Inula), Asteraceae Samphire, Marsh (Salicornia), Amaranthaceae Samphire, Rock (Crithmum), Apiaceae Samydaceae, in Salicaceae Sand apple (Diospyros), Ebenaceae Sand apple (Parinari), Chrysobalanaceae Sandalwood (Santalum), Santalaceae Sanicle (Sanicula), Apiaceae Sanseviera = Dracaena, Asparagaceae Santalaceae, Santalales Santalales, Eudicots Sapele (Entandrophragma), Meliaceae Sapindaceae, Sapindales Sapindales, Malvids Sapodilla (Manilkara), Sapotaceae Sapotaceae, Ericales Sapota/e (Manilkara), Sapotaceae Sapote, Black (Diospyros), Ebenaceae Sapote, Mamey (Pouteria), Sapotaceae Sarcolaenaceae, Malvales Sarcospermataceae, in Sapotaceae Sargentodoxaceae, in Lardizabalaceae Sarraceniaceae, Ericales Sarsaparilla (Smilax), Smilacaceae Sassafras, Lauraceae Sassafras, Blackheart (or Southern/Tasmanian)

(Atherosperma) Atherospermataceae Satinleaf (Chrysophyllum), Sapotaceae Satinwood, Ceylon or East Indian

(Chloroxylon), Rutaceae Satinwood, Jamaican or West Indian

(Zanthoxylum), Rutaceae Saururaceae, Piperales Sausage tree, usually (Kigelia), Bignoniaceae Sauvagesiaceae, in Ochnaceae Savory (Summer & Winter) (Satureja), Lamiaceae Saxifragaceae, Saxifragales Saxifrage (Saxifraga), Saxifragaceae Scabious, Caprifoliaceae Schefflera, Araliaceae Scheuchzeriaceae, Alismatales Schisandraceae, Austrobaileyales Schoepfiaceae, Santalales Schlegeliaceae, Lamiales Screw pines (Pandanus), Pandanaceae Scrophulariaceae, Lamiales Scyphostegiaceae, in Salicaceae Scytopetalaceae, in Lecythidaceae Sea buckthorn (Hippophae), Elaeagnaceae Sea heath (Frankenia), Frankeniaceae Sea holly (Acanthus), Acanthaceae Sea holly (Eryngium), Apiaceae Sea lavender (Limonium), Plumbaginaceae Sea lettuce (Scaevola), Goodeniaceae Sedge (usually Carex), Cyperaceae Selago, Scrophulariaceae Self-heal (Prunella & Stachys), Lamiaceae Semolina (Triticum), Poaceae Senecio, Asteraceae Senega root (Polygala), Polygalaceae Senna, Fabaceae Sesame (Sesamum), Pedaliaceae Setchellanthaceae, Brassicales She-oak (Casuarina & Allocasuarina), Casuarinaceae Shea (butter) (Vitellaria), Sapotaceae Shisham wood, (Dalbergia), Fabaceae Shoal grass (Halodule), Cymodoceaceae Shorea, Dipterocarpaceae Shortia, Diapensiaceae Silk-tassle bush (Garrya), Garryaceae Silk vine (Periploca), Apocynaceae Silvianthus, Carlemanniaceae Silver ash (Flindersia), Rutaceae Silverbells (Halesia), Styracaceae Simaroubaceae, Sapindales Simmondsiaceae, Caryophyllales Siparunaceae, Laurales Siphonodontaceae, in Celastraceae Sipo (Entandrophragma), Meliaceae Sisal (Agave), Asparagaceae Sissoo wood, (Dalbergia), Fabaceae Sladeniaceae, Ericales Smilacaceae, Liliales Smilax, Smilacaceae Smoketree/bush (Cotinus), Anacardiaceae Smoketree (Psorothamnus), Fabaceae Snake-root (Ageratum), Asteraceae Snake-root (Eupatorium), Asteraceae Snake-root, Indian (Rauwolfia), Apocynaceae Snake-root, Virginian (Aristolochia),

Aristolochiaceae Snake-root, Black (Cimifuga), Ranunculaceae Snake-root, Black (Sanicula), Apiaceae Snakewood (Acacia), Fabaceae Snakewood (Brosimum = Piratinera), Moraceae Snakewood (Colubrina), Rhamnaceae Snapdragon (Antirrhinum), Plantaginaceae Sneezewood (Ptaeroxylon), Rutaceae Snowberry (Symphoricarpos), Caprifoliaceae Snowberry (Chiococca), Rubiaceae Snowbush, Hawaiian (Breynia), Phyllanthaceae Snowdrop (Galanthus), Amaryllidaceae Soapberry (Sapindus), Sapindaceae Soapberry (Shepherdia), Elaeagnaceae Soapwood (Caryocar), Caryocaraceae Soapwood (Pithecellobium), Fabaceae Solanaceae, Solanales Solanales, Lamiids Soldanella, Primulaceae Sonneratiaceae, in Lythraceae Sorghum, Poaceae Sorrel (Rumex), Polygonaceae Sorrel, Wood (Oxalis), Oxalidaceae Souari nut (Caryocar), Caryocaraceae Soursop (Annona), Annonaceae Sourwood (Oxydendrum), Ericacaeae Southern beech (Nothofagus), Nothofagaceae Southernwood (Artemisia), Asteraceae Sow thistle (Sonchus), Asteraceae Soya (Glycine), Fabaceae Spanish moss (Tillandsia), Bromeliaceae Sparganiaceae, in Typhaceae Spectacle fruit (Wislizenia), Cleomaceae Sphaerosepalaceae, Malvales Sphenocleaceae, Solanales Sphenostemon, Paracryphiaceae Spiderwort (Tradescantia), Commelinaceae Spikenard (Nardostachys), Caprifoliaceae Spinach (Spinacia), Amaranthaceae Spinach, Indian/Malabar (Basella), Basellaceae Spindle tree (Euonymus), Celastraceae Spiraea, Rosaceae Spring beauty (Claytonia), Montiaceae Spurge (Euphorbia), Euphorbiaceae Squash (Cucurbita), Cucurbitaceae Squill (Scilla), Asparagaceae Stachyuraceae, Crossosomatales Stackhousiaceae, in Celastraceae Staff-tree (Celastrus), Celastraceae Staphyleaceae, Crossosomatales Star anise (Illicium), Schisandraceae Star apple (Chrysophyllum), Sapotaceae Starwort (Callitriche), Plantaginaceae Stemonaceae, Pandanales Stemonuraceae, Aquifoliales Sterculiaceae, in Malvaceae Stilbaceae, Lamiales Stinkwoods (see below, also Foetidia,

Lecythidaceae; Gyrocarpus, Hernandiaceae; Zieria, Rutaceae, etc.) Stinkwood, Cape or Black

(Ocotea), Luraceae

Stinkwood, Green (Jacksonia), Fabaceae Stinkwood, Grey (Jacksonia), Fabaceae Stinkwood, Red (Prunus), Rosaceae Stinkwood, White (Celtis), Cannabaceae Stixaceae, Brassicales Stock (Matthiola), Brassicaceae Stock, Virginia (Malcolmia), Brassicaceae Stonecrop (Sedum), Crassulaceae Storax (Styrax), Styracaceae Strawberry (Fragaria), Rosaceae Strawberry tree (Arbutus), Ericaceae Strelitziaceae, Zingiberales Streptocarpus, Gesneriaceae Stripeseed (Piriqueta), Passifloraceae Strombosiaceae, Santalales Strychnine (Strychnos), Loganiaceae Stylidiaceae, Asterales Stylobasiaceae, in Surianaceae Styracaceae, Ericales Styrax, Styracaceae Sugar cane (Saccharum), Poaceae Sumac(h) (Rhus), Anacardiaceae Sun cup (Camissonia), Onagracaeae Sundew (Drosera), Droseraceae Sunflower (Helianthus), Asteraceae Sunrose (Helianthemum), Cistaceae Surette (Byrsonima), Malpighiaceae Surianaceae, Fabales Sweet chestnut (Castanea), Fagaceae Sweet flag (Acorus), Acoraceae Sweet gale (Myrica), Myricaceae Sweet pea (Lathyrus), Fabaceae Sweet potato (Ipomoea), Convolvulaceae Sweetgum (Liquidambar), Altingiaceae Sweetleaf (Sauropus), Phyllanthaceae Sweetleaf (Stevia), Asteraceae Sweetleaf (Symplocos), Symplocaceae Sweetshrub (Calycanthus), Calycanthaceae Sweetsop (Annona), Annonaceae Sweetspire (Itea), Iteaceae Sycamore [UK] (Acer), Sapindaceae Symphoricarpus, Caprifoliaceae Symplocaceae, Ericales

Tabonuco (Dacryodes), Burseraceae Taccaceae, in Dioscoreaceae (provisionally) Talipot (Corypha), Arecaceae Tamanu oil (Calophyllum), Calophyllaceae Tamaricaceae, Caryophyllales Tamarind (Tamarindus), Fabaceae Tamarisk (Tamarix), Tamaricaceae Tambalacoque (Sideroxylon), Sapotaceae Tan oak (Lithocarpus), Fagaceae Tansy (Tanacetum), Asteraceae Tapioca (Manihot), Euphorbiaceae Tara vine (Actinidia), Actinidiaceae Tarata (Pittosporum), Pittosporaceae Taro (Colocasia, also Alocasia, Amorphophallus,

Crytosperma & Xanthosoma), Araceae Tarragon (Artemisia), Asteraceae Tasmanian laurel (Anopterus), Escalloniaceae Tawhai (Nothofagus), Nothofagaceae Tea (Camellia), Theaceae Tea tree (Melaleuca), Myrtaceae Tea, Bogata (Symplocos), Symplocaceae Tea, Labrador (Ledum), Ericaceae Teak (Tectona), Lamiaceae Teasel (Dipsacus), Caprifoliaceae Tecophilaeaceae, Asparagales Tef (Eragrostis), Poaceae Tepuianthus, Thymelaeaceae Tequila (Agave), Asparagaceae Ternstroemiaceae, in Pentaphylacaceae Tetracentraceae, in Trochodendraceae Tetrachondraceae, Lamiales Tetramelaceae, Cucurbitales Tetrameristaceae, Ericales Theaceae, Ericales Theligonaceae, in Rubiaceae Themidaceae, in Asparagaceae Theophrastaceae, in Primulaceae Thesiaceae, in Santalaceae Thismia, Burmanniaceae (provisionally) Thismiaceae, in Burmanniaceae (provisionally) Thistle (Carduus, Cirsium, Onopordum,

etc.), Asteraceae Thistle, Globe (Echinops), Asteraceae Thistle, Sow (Sonchus), Asteraceae Thomandersiacae, Lamiales Thorn apple (Datura), Solanaceae Thrift (Armeria), Plumbaginaceae Throatwort (Trachelium), Campanulaceae Thunbergia, Acanthaceae Thurniaceae, Poales Thyme (Thymus), Lamiaceae Thymelaeaceae, Malvales Ticodendraceae, Fagales Tigerwood (Astronium), Anacardiaceae Tigerwood (Coula), Coulaceae Tigerwood (Lovoa), Meliaceae Tiliaceae, in Malvaceae Tipu (Tipuana), Fabaceae Titi (Cliftonia & Cyrilla), Cyrillaceae Toadflax (Linaria), Plantaginaceae Tobacco (Nicotiana), Solanaceae Tolu balsam (Myroxylon), Fabaceae Tomato (Solanum), Solanaceae Toon (Toona), Meliaceae Toquilla straw plant (Carludovica), Cyclanthaceae Torchwood (Amyris), Rutaceae Torenia, Linderniaceae Torote (Bursera), Burseraceae Torricelliaceae, Apiales Tovariaceae, Brassicales Trapa, Lythraceae Trapaceae, in Lythraceae Travellers’ palm (Ravenala), Strelitziaceae Tree of heaven (Ailanthus), Simaroubaceae Tremandraceae, in Elaeocarpaceae Tribelaceae, in Escalloniaceae Trichostephanus, in Salicaceae (provisionally) Trigger plant (Stylidium), Stylidiaceae Trigoniaceae, Malpighiales Trillium, Melianthiaceae Trimeniaceae, Austrobaileyales Trincomali wood (Berrya), Malvaceae Triuridaceae, Panadanales Trochodendraceae, Trochodendrales Trochodendrales, Eudicots Tropaeolaceae, Brassicales Trumpet vine (Campsis), Bignoniaceae Tulip Tulipa, Liliaceae Tulip tree (Liriodendron), Magnoliaceae Tummy wood (Careya), Lecythidaceae Tung tree (Vernicia), Euphorbiaceae Tupelo (Nyssa), Cornaceae Turmeric (Curcuma), Zingiberaceae Turneraceae, in Passifloraceae Turnip (Brassica), Brassicaceae Turtle grass (Thalassia), Hydrocharitaceae Tutsan (Hypericum), Clusiaceae Typhaceae, Poales

Udo (Aralia), Araliaceae Ulmaceae, Rosales Ulmo (Eucryphia), Cunoniaceae Unicorn plant (Proboscidea), Martyniaceae Urticaceae, Rosales

Valerian (Valeriana), Caprifoliaceae Valerianaceae, in Caprifoliaceae Vanilla, Orchidaceae Velloziaceae, Pandanales Venus’ fly-trap (Dionaea), Droseraceae Verbascum, Scrophulariaceae Verbena, Verbenaceae Verbenaceae, Lamiales Veronica, Plantaginaceae Vervain (Verbena), Verbenaceae Vetch (Vicia), Fabaceae Vetchling (Lathyrus), Fabaceae Vetiver (Chrysopogon), Poaceae Viburnum, Adoxaceae Vietnamese coriander (Persicaria), Polygonaceae Violaceae, Malpighiales Violet (Viola), Violaceae Violet, African (Saintpaulia), Gesneriaceae Viscaceae, Santalaceae Vitaceae, Eudicots Vivianiaceae, Geraniales Vochysiaceae, Myrtales

Wallflower (Erysimum & Cheiranthus), Brassicaceae Walnut (Juglans), Juglandaceae Walnut, African (Coula), Coulaceae Water caltrop (Trapa), Lythraceae Water canna (Thalia), Marantaceae Water chestnut (Eleocharis), Cyperaceae Water hawthorn (Aponogeton), Aponogetonaceae Water hemlock (Cicuta), Apiaceae Water hyacinth (Eichhornia), Pontederiaceae Water lily, various genera in Nymphaeaceae

& Nelumbonaceae & Cabombaceae Water melon (Citrullus), Cucurbitaceae Water milfoil (Myriophyllum), Haloragaceae Water pennywort (Hydrocotyle), Araliaceae Water plantain (Alisma), Alismataceae Water plantain, Floating (Luronium), Alismataceae Water poppy (Hydrocleys), Alismataceae Water-shield (Brasenia), Cabombaceae Water-shield, Carolina (Cabomba), Cabombaceae Water starwort (Callitriche), Plantaginaceae Watercress (Nasturtium), Brassicaceae Waterleaf (Hydrophyllum), Boraginaceae Waterleaf (Talinum), Talinaceae Waterwort (Elatine), Elatinaceae Wattle (Acacia), Fabaceae Wax myrtle (Myrica), Myricaceae Waxplant (Hoya), Apocynaceae Weigela, Caprifoliaceae Weld (Reseda), Resedaceae Wellstediaceae, in Boraginaceae West Indian bay tree (Pimenta), Myrtaceae West Indian boxwood (Casearia), Salicaceae West Indian boxwood (Phyllostylon), Ulmaceae West Indian boxwood (Tabebuia), Bignoniaceae Wheat (Triticum), Poaceae White beech (Gmelina), Lamiaceae White mangrove (Avicenna), Acanthaceae White mangrove (Conocarpus, Languncularia

& Lumnitzera), Combretaceae White pear, African (Apodytes), Icacinaceae (provisionally) Widgeon grass (Ruppia), Ruppiaceae Willow (Salix), Salicaceae Willow myrtle (Agonis), Myrtaceae Willowherb (Epilobium), Onagraceae Wingnut (Pterocarya), Juglandaceae Winter cress (Barbarea), Brassicaceae Winter’s bark (Drimys), Winteraceae Winteraceae, Canellales Winterberry (Ilex), Aquifoliaceae Wintergreen (Gaultheria), Ericaceae Wintergreen (Pyrola and related genera), Ericaceae Wintergreen oil, various familes and genera, including above, also Betula & Spiraea Wire-netting bush (Corokia), Argophyllaceae Wishbone flower (Torenia), Linderniaceae Witch hazel (Hamamelis), Hamamelidaceae Wolfberry (Lycium), Solanacaeae Wolfberry (Symphoricarpos), Caprifoliaceae Wood sage (Teucrium), Lamiaceae Wood sorrel (Oxalis), Oxalidaceae Woodrush (Luzula), Juncaceae Wormwood (Artemisia), Asteraceae Woundwort (Stachys), Lamiaceae

Xanthophyllaceae, in Polygalaceae Xanthorrhoeaceae, Asparagales Xeronemataceae, Asparagales Ximeniaceae, Santalales Xyridaceae, Poales

Yam (Dioscorea), Dioscoreaceae Yam [mainly USA] (Ipomoea), Convolvulaceae Yarrow (Achillea), Asteraceae Yellow bird’s nest (Monotropa), Ericaceae Yellow-eyed grass (Sisyrinchium), Iridaceae Yellow-eyed grass (Xyris), Xyridaceae Yerba mansa (Anemopsis), Saururaceae Yerba mate (Ilex), Aquifoliaceae Yerba santa (Eriodictyon), Boragninaceae Yerba santa (Piper), Piperaceae Ylang-ylang (Cananga), Annonaceae Yucca, Asparagaceae

Zannichelliaceae, in Potamogetonaceae Zapote = Sapote (Manilkara), Sapotaceae Zebra-wood, usually Microberlinia, Fabaceae,

also Connarus, Connaraceae and others Zelkova, Ulmaceae Zingiberaceae, Zingiberales Zingiberales, Monocots Zinnia, Asteraceae Zosteraceae, Alismatales Zucchini (Cucurbita), Cucurbitaceae Zygophyllaceae, Zygophyllales Zygophyllales, Fabids